OpenStudy (anonymous):

a 1 2 - + - + - 6 4 6

OpenStudy (mayankdevnani):

what is this ?

OpenStudy (anonymous):

they are fractions

OpenStudy (mayankdevnani):

oh!!!

OpenStudy (anonymous):

add these fractions together

OpenStudy (mayankdevnani):

\[\large \bf \frac{a}{6}+\frac{1}{4}+\frac{2}{6}\]

OpenStudy (mayankdevnani):

is this your question?

OpenStudy (anonymous):

yes exactly sorry mine was confusing

OpenStudy (mayankdevnani):

use LATEX

OpenStudy (anonymous):

what?

OpenStudy (mayankdevnani):

to type in fraction form"- Type this in equation editor :- then type this :- `frac{x}{y}` =\(\frac{x}{y}\)

OpenStudy (mayankdevnani):

understood? @mef1717

OpenStudy (mayankdevnani):

any way,here is solution :- First take LCM .

OpenStudy (anonymous):

from the denominator?

OpenStudy (mayankdevnani):

yeah

OpenStudy (princeharryyy):

(2+a)/6 + 1/4 =>(8+4a +6)/24 = (14+4a)/24 @mef1717

OpenStudy (mayankdevnani):

LCM is always taken from DENOMINATOR

OpenStudy (mayankdevnani):

then what you get after taking LCM

OpenStudy (anonymous):

thank you @princeharryyy

OpenStudy (anonymous):

okay so we take the lcm from the denominator and it's 24?

OpenStudy (mayankdevnani):

do you understand what @princeharryyy answered?

OpenStudy (anonymous):

kind like is the denominator 24 because its the lcm?

OpenStudy (mayankdevnani):

okk..24 is the LCM

OpenStudy (anonymous):

okay then what?

OpenStudy (mayankdevnani):

then, we get, \[\large \bf \frac{4a+6+8}{24}=\frac{4a+14}{24}=\frac{4a}{24}+\frac{14}{24}=\frac{a}{6}+\frac{7}{12}\]

OpenStudy (mayankdevnani):

so \[\large \bf \frac{a}{6}+\frac{7}{12}=\color{red}{Answer}\]

OpenStudy (anonymous):

so we have to make the denominators the same?

OpenStudy (mayankdevnani):

no,this is final answer because we break the common denominator from the answer

OpenStudy (mayankdevnani):

from priche have answered

OpenStudy (mayankdevnani):

*prince

OpenStudy (princeharryyy):

yes!!! @mef1717 while adding you should have the denominators same

OpenStudy (anonymous):

OHH thank you to you both!

OpenStudy (mayankdevnani):

welcome :)

OpenStudy (princeharryyy):

Gud Luck !!! hope u'll do the rest byurself @mef1717 ... @mayankdevnani explained it to u beautifully ... n u r brilliant as well ...

OpenStudy (mayankdevnani):

and @princeharryyy read out this :- http://openstudy.com/code-of-conduct In short,i want to say that don't give DIRECT ANSWERS otherwise you will be suspended from OPEN STUDY :) ----->ADVISE FROM me

OpenStudy (princeharryyy):

I wrote each step @mayankdevnani ... I think

OpenStudy (princeharryyy):

direct answes mean only the final answers @mayankdevnani

OpenStudy (princeharryyy):

answers*

OpenStudy (mayankdevnani):

yeah...thats the reason..this is somewhat like DIRECT ANSWER.. I mean that first give hint,just like me example :- take LCM first. like that and then if he reply i don't understand then tell him whole answer with each and every step beautifully :) Hope you understand. @princeharryyy

OpenStudy (princeharryyy):

ohk Brother !!! n thanks for telling ... well! I have my xam tomorrow I should leave ....

OpenStudy (mayankdevnani):

BEST OF LUCK for your EXAM :)

OpenStudy (princeharryyy):

tnku ... @mayankdevnani

OpenStudy (mayankdevnani):

welcome :)